OpenStudy (anonymous):

What is the square root of 75?

OpenStudy (hellokitty17):

8.66025

OpenStudy (johnweldon1993):

Hint* how else can you write 75 using the factors of 75?

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

ITS BASICALLY 75X75, SO TRY THAT

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm L}\left(\sqrt{75}\right)=\sqrt{81}+\frac{1}{2\sqrt{81}}\left(75-81\right) =9-\frac{6}{18}=8.\bar{6} }\)

OpenStudy (solomonzelman):

approximation, using linearization

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

8.66025 multiplied by itself equals 74. 99993006, so do not use that as your answer.

OpenStudy (anonymous):

ok thanks for the advice! :)

OpenStudy (anonymous):

Of course!

OpenStudy (anonymous):

Solomon can u help me after ur done?

OpenStudy (solomonzelman):

8.66025 is an exact answer (from calculator that uses an (Infinity)th degree polynomial approximation, well not really infinity, but you know...)